OpenStudy (barrelracer011):

HELP ASAP!! ALGEBRA 1!! WILL MEDAL!! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (barrelracer011):

A. 12x + 4y = -12 15x = -30 B. 8x + 4y = -4 14x = -10 C. 6x - 8y = -36 -6y = -42 D. 6x - y = -15 3y = 9

OpenStudy (barrelracer011):

@Nateyboo can you help?

OpenStudy (anonymous):

Yea ill try. Give me a minute (:

OpenStudy (anonymous):

I would suggest first solving the system of equations for x and y...

OpenStudy (anonymous):

its A in my opinion

OpenStudy (anonymous):

The solutions to your equations (6x + 2y = -6 & 3x - 4y = -18) are: x= -2 and y= 3

OpenStudy (barrelracer011):

so you would plug in X=-2 and y=3 to each equation and see which one is true?

OpenStudy (anonymous):

12x + 4y = -12 15x = -30 15x = -30 x = -30/15 = -2 12(-2) + 4y = -12 -24 + 4y = -12 4y = -12 + 24 4y = 12 y = 12/4 = 3

OpenStudy (anonymous):

so A is right option

OpenStudy (barrelracer011):

Ok thank you so much!

OpenStudy (anonymous):

Welcome deary :)

OpenStudy (anonymous):

definitely A.

OpenStudy (barrelracer011):

haha thanks @Nateyboo

OpenStudy (anonymous):

I know I was late . lol But Your welcome anyways (:

OpenStudy (anonymous):

You did a great job Natey no doubt :)

OpenStudy (barrelracer011):

Your fine its great that I have someone confirming the answer.

OpenStudy (anonymous):

Thanks . :D

OpenStudy (barrelracer011):

Your welcome!